### v3

@@ Abstract
This paper introduces a computationally efficient method that globally converges to B-stationary points of mathematical programs with equilibrium constraints (MPECs) in a finite number of iterations. B-stationarity is necessary for optimality and means that no feasible first-order direction can improve the objective. Given a feasible point of an MPEC, B-stationarity can be certified by solving a linear program with equilibrium constraints (LPEC) constructed at this point. The proposed method solves a sequence of LPECs, which either certify B-stationarity or provide an active-set estimate for the complementarity constraints, along with nonlinear programs (NLPs) – referred to as branch NLPs (BNLPs) – obtained by fixing the active set in the MPEC. A BNLP is more regular than the original MPEC, easier to solve, and with the correct active set, its solution coincides with that of the MPEC. We show that, unless the current iterate is B-stationary, these combinatorial LPECs need not be solved to optimality; for convergence, it suffices to compute a nonzero feasible point, yielding significant computational savings. The method proceeds in two phases: the first identifies a feasible BNLP or a stationary point of a constraint infeasibility minimization problem, and the second solves a finite sequence of BNLPs until a B-stationary point of the MPECis found. We established finite convergence under the MPEC-MFCQ. Numerical experiments and an open-source software implementation show that the proposed method is more robust and faster than relaxation-based and mixed-integer NLP approaches, even on medium to large-scale instances.
@@ 1 Introduction
We highlight key differences from other hybrid approaches. Unlike Lin and Fukushima Lin2006 , our method solves LPECs for active set guesses rather than using multipliers, eliminating MPEC-LICQ requirements and allowing convergence to B-stationary points that are not S-stationary. Unlike Kazi et al. Kazi2024 , we never apply LPEC steps directly, ensuring all Phase II iterates remain feasible and guaranteeing finite termination.
@@
Contributions: This paper presents a new method, called MPEC optimizer (MPECopt) that computes B-stationary points by solving a finite number of LPECs and NLPs. In our method, the LPECs additionally have a trust-region constraint on the step lengths. We discuss in detail two possible different LPEC formulations and their relation.
We show that, unless we need to certify B-stationarity, it is sufficient to find a feasible (improving) point of the LPEC to ensure progress. Moreover, we discuss special cases when LPECs are easy to solve, and cases where a MILPs reformulation can be solved efficiently.
We prove that, by finding a feasible point of an LPEC constructed at a point sufficiently close to the feasible set of the MPEC, the active set obtained from the LPEC corresponds to a BNLP that is guaranteed to be feasible. In practice, this helps to find feasible BNLPs in Phase I by combining well-established regularization methods and an LPEC solver. Phase II iterates over the feasible BNLPs while strictly decreasing the objective until a B-stationary point is found. We prove that, under the MPEC-MFCQ, Phase II finds a B-stationary point in a finite number of steps. Furthermore, we compare the new method with several relaxation-based methods, and a Mixed-Integer Nonlinear Programming (MINLP) reformulation, on the MacMPEC and a synthetic large-scale nonlinear MPEC benchmark. Higher success rates and generally faster computation times are reported, and solving LPECs turns out not to be a computational bottleneck, even for large problem instances. An open-source implementation of the method is provided at github.com/nosnoc/mpecopt.
@@
Outline: Section 2 recalls some preliminaries. Section 3 provides a detailed statement of MPECopt, discusses how to formulate and solve LPECs. Section 4 gives a detailed convergence analysis of the new method. Section 5 provides all implementation details and numerical results on two benchmark test sets. Finally, Section 6 provides concluding remarks.
@@ 2 Preliminaries
Similarly to $x=(x_{0},x_{1},x_{2})$, the vector $d$ is partitioned as $d=(d_{0},d_{1},d_{2})\in\mathbb{R}^{p}\times\mathbb{R}^{m}\times\mathbb{R}^{m}$. Observe that at a feasible point $x\in\Omega$ all constraint residuals above are zero but they are kept for convenience. In contrast to the classical linearized feasible cone, here the combinatorial structure is kept for the degenerate index set ${\mathcal{I}}_{00}(x)$ and $\mathcal{F}_{\Omega}(x)$ is nonconvex, if $\mathcal{I}_{00}$ is nonempty. It is a union of a finite number of convex polyhedral cones.
@@
A point $x^{*}$ is called B-stationary if $d=0$ solves the following LPEC:
@@
Let us define multiplier-based stationary concepts for MPECs and some auxiliary NLPs associated with the MPEC (1). Some of these problems will be solved by the algorithm introduced in this paper, while others serve to define additional concepts relevant to MPECs. To this end, we introduce further index sets.
Let $x^{*}\in\Omega$ and ${\mathcal{D}}_{1}(x^{*}),{\mathcal{D}}_{2}(x^{*})$ be a partition of the biactive set ${\mathcal{I}}_{00}(x^{*})$, with $\mathcal{D}_{1}(x^{*})\cup\mathcal{D}_{2}(x^{*})={\mathcal{I}}_{00}(x^{*})$ and $\mathcal{D}_{1}(x^{*})\cap\mathcal{D}_{2}(x^{*})=\emptyset$.
@@
It can be seen that the following holds for $x^{*}\in\Omega$ Scheel2000 :
@@
with the MPEC Lagrange multipliers $\mu\in{\mathbb{R}}^{n_{c}}$, $\nu\in{\mathbb{R}}^{m}$ and $\xi\in{\mathbb{R}}^{m}$. By applying the KKT conditions to the TNLP (9), we obtain W-stationarity concept, and applying them to the RNLP yields S-stationary Scheel2000 . Further restrictions on the multipliers $\nu^{*}_{i},\xi^{*}_{i},\ i\in{\mathcal{I}}_{00}(x^{*})$, lead to M-, C-, and A-stationarity, which are not directly related to the KKT conditions of an auxiliary NLPs, see e.g.  Kim2020 ; Outrata1999 ; Scheel2000 . These concepts are summarized in the next definition.
@@
We note that if ${\mathcal{I}}_{00}=\emptyset$, then all stationarity concepts are identical and collapse to the concept of S-stationarity. From the definitions it follows that S-stationarity is the most restrictive since it has the tightest condition in the multipliers $\nu^{*}_{i},\xi^{*}_{i},i\in{\mathcal{I}}_{00}(x^{*})$. Weaker stationarity concepts, ordered by strength, are M, C, A, and W, with the stronger ones implying the weaker ones. The relationship between multiplier-based stationarity (Def. 2.3) and B-stationarity (Def. 2.1) is established in the following theorem.
@@
In summary, without the arguably restrictive MPEC-LICQ, B-stationarity cannot be verified using multiplier-based concepts. This assumption is essential, as the next weaker condition, MPEC-MFCQ, does not guarantee S-stationarity (cf. (Scheel2000, , Example 3)). Weaker termination criteria are unsatisfactory, since even M-stationarity allows trivial descent directions Leyffer2007 . Even under MPEC-LICQ, regularization Ralph2004 ; Hoheisel2013 ; Kanzow2015 and some active set methods (Kirches2022, , Section 3) may converge to points weaker than S-stationary. Therefore, we propose a method using the LPEC (6), to correctly certify B-stationarity.
@@ 3.1 Summary of the approach
1: Call Phase I to find a feasible point $x^{0}$ of the MPEC (1)
@@
If a feasible point is found in Phase I, Phase II creates a finite sequence of iterates $x^{k}\in\Omega$ by solving BNLPs and LPECs. If Phase II finds that any BNLP is unbounded, then the MPEC is unbounded, and the algorithm terminates. The LPECs are similar to (6), but they are defined at the linearization point $x^{k}$ and include an additional trust-region constraint $\|d\|_{\infty}\leq\rho$, $\rho>0$. These LPECs are compactly denoted by $\mathrm{LPEC}(x,\rho)$; see Section 3.2 for more details.
At each major iteration $k$, an inner loop with counter $l$ solves $\mathrm{LPEC}(x^{k},\rho^{k,l})$ for $d^{k,l}$ until either a feasible point $x^{k+1}=x^{k,l}$ with an improved objective value is found (lines 7–10), or the algorithm terminates due to $d^{k,l}=0$ (line 6), or unboundedness (line 8).
@@ 3.2.1 On different LPEC formulations
Next, we show that the assumption of MPEC-ACQ (i.e., that ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$), cannot be relaxed if one wishes to verify geometric B-stationarity (cf. Theorem 2.1) by solving the LPEC 12, which verifies algebraic/linearized B-stationarity (cf. Def. 2.1).
@@
The LPEC (12) is well-defined at feasible points. However, in the Phase I algorithm of Section 3.3, we must also consider LPECs at infeasible points, where defining the index sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+},{\mathcal{I}}_{00}$ is unclear. Existing strategies for identifying active sets from infeasible points often require overly restrictive assumptions Lin2006 . Inspired by (Scholtes2004, , Sec. 3.2), alternative approaches instead include all linearized constraints, yielding the LPEC Kirches2022 ; Leyffer2007 :
@@
$\begin{split}0<\rho<\bar{\rho}=\min\Big\{&\{{x^{k}_{1,i}}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{{x^{k}_{2,i}}\mid i\in{\mathcal{I}}_{0+}^{k}\}\\ &\cup\{\frac{c_{i}(x^{k})}{\|\nabla c_{i}(x^{k})\|}\mid i\in\bar{{\mathcal{A}}}^{k},\|\nabla c_{i}(x^{k})\|\neq 0\}\Big\},\end{split}$ (16)
@@
In simpler terms, at a feasible point $x^{k}$ and for a sufficiently small trust-region radius, all branches that are not biactive at $x^{k}$ are eliminated, and the linearizations of inactive constraints remain inactive. Thus, the feasible sets of the two LPECs coincide. On the other hand, for $\rho\geq\bar{\rho}$, the global minimizer $d=0$ of (12) may correspond only to a local minimizer of (15), cf. Example 3.2. Hence, global optimality of $d=0$ is sufficient but not necessary to verify B-stationarity via (15), cf. Scholtes2004 .
[…]
@@
Let $x^{k}\in\Omega$ be a feasible point of the MPEC (1) and suppose that the MPEC-ACQ holds at this point. Let $d$ be any feasible point of the $\mathrm{LPEC}(x^{k},\rho)$ in (15) for $\rho\in(0,\bar{\rho})$, where $\bar{\rho}$ is defined in (16). Then $x^{k}$ is a feasible point of $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d),{\mathcal{I}}_{2}(x^{k}+d))$.
@@ 3.2.2 On solving LPECs
We summarize several favorable cases that frequently arise. First, if the LPEC contains only bound and complementarity constraints, it can be solved in linear time (Kirches2022, , Propositions 2.3 and 2.4). This algorithm is incorporated into our implementation. Second, at a feasible point $x^{k}$, if ${\mathcal{I}}_{00}(x^{k})=\emptyset$, then the LPEC (12) reduces to a standard LP. According to (10), the LPEC can then be solved by enumerating all branch LPs, which is not computationally expensive when $|{\mathcal{I}}_{00}(x^{k})|$ is small. Third, if $x^{k}$ is not B-stationary, it is unnecessary to solve all branch LPs; it suffices to find a feasible direction $d$ with $\nabla f(x^{k})^{\top}d<0$ (cf. Corollary 4.2). Moreover, in the LPECs arising in Phase I of our method, it is sufficient to simply find a feasible point (cf. Theorem 4.1). In a MILP formulation, the solver can therefore be terminated as soon as such a feasible point is obtained. Section 5 shows that this results in a significant reduction in the number of LPs that need to be solved in the MILP algorithm. Finally, if $x^{k}$ is S-stationary, B-stationarity can be verified by solving a single LP (cf. Proposition 3.1). In summary, except when verifying B-stationarity, it is rarely necessary to solve the LPEC to global optimality – and even then, the task often reduces to solving a single LP.
In this paper, we focus on arguably the easiest-to-implement approach for LPECs, and which does not suffer from the drawbacks of regularization methods. We consider the equivalent MILP formulation of LPEC (15Hu2012 :
@@
The MILP (18) can be solved with state-of-the-art commercial and open-source solvers such as Gurobi Gurobi or HiGHS Huangfu2018 . We discuss some special properties of the MILP (18), which are favorable for MILP solvers Gurobi ; Huangfu2018 .
We denote the optimal objective value of (18) by $V_{\mathrm{MILP}}(x^{k})$, and of its LP relaxation, that is, a problem where $y\in\{0,1\}^{m}$ is replaced by $y\in[0,1]^{m}$, by $V_{\mathrm{LP}}(x^{k})$. Thus, it holds that $V_{\mathrm{LP}}(x^{k})\leq V_{\mathrm{MILP}}(x^{k})$. The tighter the relaxation, i.e., the smaller $V_{\mathrm{MILP}}(x^{k})-V_{\mathrm{LP}}(x^{k})$ is, the less branching may be needed in MILP solvers Wolsey1998 . Big-M formulations can be inefficient because they can lead to rather loose relaxations of the MILP Hu2008 . However, for $\rho<\frac{M}{2}$, the trust-region constraint (18e) always dominates the bounds on $d$ from the big-M constraints (18c) and (18d). Thus, $M$ only needs to be large enough to ensure the feasibility of (18c)–(18d). A suitable choice is $M=\max(x_{1}^{k},x_{2}^{k})+\rho$. In addition, observe that the constraint (18e) makes the feasible set of the MILP compact. Moreover, for feasible points $x^{k}$, the trust-region radius $\rho$ can be arbitrarily small, leading to a smaller $V_{\mathrm{LP}}(x^{k})$ and thus tighter relaxations of the MILP.
Furthermore, given a feasible complementarity pair, e.g. from the previous BNLP, a feasible binary vector $y$ can be constructed, providing an upper bound for $V_{\mathrm{MILP}}(x^{k})$. The availability of an upper bound can reduce branching in branch-and-bound MILP algorithms Wolsey1998 , and our implementation includes the procedure for computing such a feasible binary vector.
Our numerical experiments indicate that these factors typically result in not large numbers of LP that need to be solved when solving the MILP (18). Interestingly, in our experiments the MILP formulations for LPECs are often substantially faster than solving the LPEC (15) with noncombinatorial regularization-based methods. We discuss a special case that frequently arises in our numerical experiments when $x^{*}\in\Omega$ is an S-stationary point. In this case, the MILPs are particularly easy to solve since the relaxation is tight, i.e., $V_{\mathrm{MILP}}(x^{*})=V_{\mathrm{LP}}(x^{*})$. Thus, verifying B-stationarity reduces to solving a single LP. We prove this for the reduced LPEC (12). However, if $\rho<\bar{\rho}$ (with $\bar{\rho}$ from (16)) and the preprocessing step for fixing integers is applied, then the feasible sets of the full MILP (18) and reduced MILP (19) coincide.
@@
Let $x^{*}\in\Omega$ be an S-stationary point of the MPEC (1) at which the MPEC-ACQ holds. Consider the MILP reformulation of the reduced LPEC (12):
@@
While Algorithm 1 solves NLPs and MILPs similar to many MINLP methods Belotti2013 , there are significant differences in termination criteria and the particular subproblems that are solved. Nonlinear branch-and-bound algorithms terminate when the optimality gap falls below a threshold Belotti2013 , which does not directly relate to B-stationarity unless the global optimum is found. Outer approximation algorithms also solve sequences of NLPs and MILPs Belotti2013 , but their MILPs include linearized constraints from multiple different linearization points, making them different from the MILP (18). Consequently, it is not immediately clear when these methods terminate at B-stationary points. In summary, Algorithm 1 is not a MINLP algorithm; rather, it employs MILP solvers as a computational tool for solving LPECs. It aims to find a locally stationary point of an MPEC, which—unlike in MINLPs – is a well-defined notion.

### v4

@@ Abstract
This paper introduces a computationally efficient method that converges globally to B-stationary points of mathematical programs with equilibrium constraints (MPECs). B-stationarity is necessary for optimality and means that no feasible first-order direction can improve the objective. It can be certified by solving a linear program with equilibrium constraints (LPEC) constructed at a given feasible point. The proposed method solves a finite sequence of LPECs, which either certify B-stationarity or provide an active-set estimate for the complementarity constraints, along with branch nonlinear programs (BNLPs) obtained by fixing the complementarity active set in the MPEC. In particular, the method proceeds in two phases: the first identifies a feasible BNLP or a stationary point of a constraint infeasibility minimization problem, and the second solves a sequence of BNLPs until a B-stationary point of the MPEC is found. We prove that under the MPEC-MFCQ, the method requires solving only a finite number of BNLPs and LPECs for convergence. Moreover, we show that, unless the current iterate is B-stationary, the combinatorial LPECs need not be solved to optimality. For convergence, it suffices to compute a nonzero feasible point, which in practice often requires solving a single linear program, yielding significant computational savings. Numerical experiments show that the proposed method is more robust and faster than relaxation-based methods and mixed-integer NLP reformulations (which, in contrast to the proposed approach, do not provide a certificate of B-stationarity), even on medium- to large-scale instances.
@@ 1 Introduction
We highlight key differences from other hybrid approaches. Unlike Lin and Fukushima Lin2006 , our method solves LPECs for active set guesses rather than using multipliers, eliminating MPEC-LICQ requirements and allowing convergence to B-stationary points that are not S-stationary. Unlike Kazi et al. Kazi2024 , we never apply LPEC steps directly, ensuring all Phase II iterates remain feasible.
@@
Contributions: This paper presents a new method, called MPEC optimizer (MPECopt), that computes B-stationary points by solving a finite number of LPECs and NLPs in two phases. We prove that, by finding a feasible point of an LPEC constructed at a point sufficiently close to the feasible set of the MPEC, the active set obtained from the LPEC corresponds to a BNLP that is guaranteed to be feasible. For our implementation, this result guarantees that a Phase I, which combines well-established regularization methods and an LPEC solver, finds feasible points of a BNLP (in fact, already stationary points of the MPEC). Phase II either certifies B-stationarity of this point or iterates over a sequence of feasible BNLPs while strictly decreasing the objective until a B-stationary point is found. We provide a detailed convergence analysis and prove that, under the MPEC-MFCQ, Phase II is guaranteed to find a B-stationary point in a finite number of LPEC and BNLP solves.
Furthermore, we discuss in detail two different LPEC formulations, their relation, and theoretical properties. We show that, unless we need to certify B-stationarity, it is sufficient to find a feasible (improving) point of the LPEC to ensure progress, which often reduces the computational effort significantly. In practice, solving a single linear program is often sufficient. We also discuss in detail why LPECs can be solved efficiently via their MILP reformulation.
We compare the proposed method with several relaxation-based methods and a mixed-integer nonlinear programming (MINLP) reformulation on two benchmarks: the MacMPEC set and a synthetic large-scale nonlinear MPEC benchmark. Higher success rates and generally faster computation times are reported. Interestingly, solving LPECs turns out not to be a computational bottleneck in practice, even for large problem instances. We highlight that most other methods with an available implementation – including those we compare to – do not, in general, provide a certificate of B-stationarity for generic MPECsand they may converge to a point with a first-order descent direction. Additionally, an open-source implementation of the method is provided at github.com/nosnoc/mpecopt.
@@
Outline: Section 2 recalls some preliminaries. Section 3 provides a detailed statement of MPECopt, discusses how to formulate and solve LPECs. Section 4 gives a detailed convergence analysis of the new method. Section 5 provides all implementation details and numerical results on two benchmark test sets. Section 6 provides concluding remarks.
@@ 2 Preliminaries
Similarly to $x=(x_{0},x_{1},x_{2})$, the vector $d$ is partitioned as $d=(d_{0},d_{1},d_{2})\in\mathbb{R}^{p}\times\mathbb{R}^{m}\times\mathbb{R}^{m}$. Observe that at a feasible point $x\in\Omega$ all constraint residuals above are zero but they are kept for convenience. In contrast to the usual linearized feasible cone, here the combinatorial structure is kept for the degenerate index set ${\mathcal{I}}_{00}(x)$ and $\mathcal{F}_{\Omega}(x)$ is nonconvex, if $\mathcal{I}_{00}$ is nonempty.
@@
A point $x^{*}$ is called (algebraic) B-stationary if $d=0$ solves the following LPEC:
@@
Next, we review multiplier-based stationary concepts for MPECs and some auxiliary NLPs associated with the MPEC (1). One of these problems will be solved by the algorithm introduced below, while others serve to define additional concepts relevant to MPECs. To this end, we introduce further index sets. Let $x^{*}\in\Omega$ and ${\mathcal{D}}_{1}(x^{*}),{\mathcal{D}}_{2}(x^{*})$ be a partition of the biactive set ${\mathcal{I}}_{00}(x^{*})$, with $\mathcal{D}_{1}(x^{*})\cup\mathcal{D}_{2}(x^{*})={\mathcal{I}}_{00}(x^{*})$ and $\mathcal{D}_{1}(x^{*})\cap\mathcal{D}_{2}(x^{*})=\emptyset$.
@@
It can be seen that the following holds locally around a point $x^{*}\in\Omega$ Scheel2000 :
@@
with the MPEC Lagrange multipliers $\mu\in{\mathbb{R}}^{n_{c}}$, $\nu\in{\mathbb{R}}^{m}$ and $\xi\in{\mathbb{R}}^{m}$.
By applying the KKT conditions to the TNLP (9), we obtain W-stationarity concept, and applying them to the RNLP yields S-stationary Scheel2000 . Further restrictions on the multipliers $\nu^{*}_{i},\xi^{*}_{i},\ i\in{\mathcal{I}}_{00}(x^{*})$, lead to M-, C-, and A-stationarity, which are not directly related to the KKT conditions of an auxiliary NLPs, see e.g.  Kim2020 ; Outrata1999 ; Scheel2000 . These concepts are summarized in the next definition.
@@
Note that, if ${\mathcal{I}}_{00}=\emptyset$, then all stationarity concepts are identical and collapse to the concept of S-stationarity. From the definitions it follows that S-stationarity is the most restrictive since it has the tightest condition in the multipliers $\nu^{*}_{i},\xi^{*}_{i},i\in{\mathcal{I}}_{00}(x^{*})$. Weaker stationarity concepts, ordered by strength, are M, C, A, and W, with the stronger ones implying the weaker ones.
The relationship between multiplier-based stationarity (Def. 2.3) and B-stationarity (Def. 2.1) is established in the following theorem.
@@
In summary, without assuming MPEC-LICQ, B-stationarity cannot be verified using multiplier-based concepts. This assumption is essential, as the next weaker condition, MPEC-MFCQ, does not guarantee S-stationarity (cf. (Scheel2000, , Example 3)). Weaker termination criteria than B-stationarity may be considered unsatisfactory, since even M-stationarity allows first-order descent directions Leyffer2007 . Even under MPEC-LICQ, regularization Ralph2004 ; Hoheisel2013 ; Kanzow2015 and some active set methods (Kirches2022, , Section 3) may converge to points weaker than S-stationary. Therefore, we propose a method using the LPEC (6), to certify B-stationarity.
@@ 3.1 Summary of the approach
1: Call Phase I to find a feasible point $x^{0}$ of the MPEC (1); set $k=0$
@@
If a feasible point is found in Phase I, Phase II creates a finite sequence of iterates $x^{k}\in\Omega$ by solving BNLPs and LPECs. If Phase II finds that any BNLP is unbounded, then the MPEC is unbounded, and the algorithm terminates. The LPECs are similar to (6), but they are defined at the linearization point $x^{k}$ and include an additional trust-region constraint $\|d\|_{\infty}\leq\rho$, $\rho>0$. These LPECs are compactly denoted by $\mathrm{LPEC}(x,\rho)$; see Section 3.2 for details.
At each major iteration $k$, an inner loop with counter $l$ solves $\mathrm{LPEC}(x^{k},\rho^{k,l})$ for $d^{k,l}$ and, possibly, a corresponding $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$, until either a feasible point $x^{k+1}=x^{k,l}$ with an improved objective value is found (lines 7–10), or the algorithm terminates due to $d^{k,l}=0$ (line 6) or the MPEC is found to be unbounded (line 8).
@@ 3.2.1 On different LPEC formulations
Next, since algebraic B-stationarity is stronger than geometric B-stationarity, we show that the assumption of MPEC-ACQ (i.e., that ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$), cannot be relaxed if one wishes to verify geometric B-stationarity (cf. Theorem 2.1) by solving the LPEC (12), which verifies algebraic B-stationarity (cf. Def. 2.1).
@@
The LPEC (12) is well-defined at feasible points. However, in the Phase I algorithm of Section 3.3, we must also consider LPECs at infeasible points, where index sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+},{\mathcal{I}}_{00}$ are note defined. Existing strategies for identifying active sets from infeasible points often require overly restrictive assumptions Lin2006 . Inspired by (Scholtes2004, , Sec. 3.2), alternative approaches instead include all linearized constraints, yielding the LPEC Kirches2022 ; Leyffer2007 :
@@
$\begin{split}0<\rho<\bar{\rho}=\min\Big\{&\{{x^{k}_{1,i}}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{{x^{k}_{2,i}}\mid i\in{\mathcal{I}}_{0+}^{k}\}\\ &\cup\{{c_{i}(x^{k})}\|\nabla c_{i}(x^{k})\|^{-1}\mid i\in\bar{{\mathcal{A}}}^{k},\|\nabla c_{i}(x^{k})\|\neq 0\}\Big\},\end{split}$ (16)
@@
In other words, at a feasible point $x^{k}$ and for a sufficiently small trust-region radius, all branches that are not biactive at $x^{k}$ are eliminated, and the linearizations of inactive constraints remain inactive. Thus, the feasible sets of the two LPECs coincide. On the other hand, for $\rho\geq\bar{\rho}$, the global minimizer $d=0$ of (12) may correspond only to a local minimizer of (15), cf. Example 3.2. Hence, global optimality of $d=0$ is sufficient but not necessary to verify B-stationarity via (15Scholtes2004 .
@@
Let $x^{k}\in\Omega$ be a feasible point of the MPEC (1) and suppose that the MPEC-MFCQ holds at this point. Let $d$ be any feasible point of the $\mathrm{LPEC}(x^{k},\rho)$ in (15) for $\rho\in(0,\bar{\rho})$, where $\bar{\rho}$ is defined in (16). Then $x^{k}$ is a feasible point of $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d),{\mathcal{I}}_{2}(x^{k}+d))$.
@@ 3.2.2 On solving LPECs
We summarize several favorable cases that frequently arise. First, if the LPEC contains only bound and complementarity constraints, it can be solved in linear time (Kirches2022, , Propositions 2.3 and 2.4). This algorithm is incorporated into our implementation. Second, at a feasible point $x^{k}$, if ${\mathcal{I}}_{00}(x^{k})=\emptyset$, then the LPEC (12) reduces to a standard LP. According to (10), the LPEC can then be solved by enumerating all branch LPs, which is not computationally expensive when $|{\mathcal{I}}_{00}(x^{k})|$ is small. Third, if $x^{k}$ is not B-stationary, it is unnecessary to solve all branch LPs; it suffices to find a feasible direction $d$ with $\nabla f(x^{k})^{\top}d<0$ (cf. Corollary 4.2). Moreover, in the LPECs arising in Phase I of our method, it is sufficient to simply find a feasible point (cf. Theorem 4.1). In a MILP formulation, the solver can therefore be terminated as soon as such a feasible point is obtained. Section 5 shows that this results in a significant reduction in the number of LPs solved in the MILP algorithm. Finally, if $x^{k}$ is S-stationary, B-stationarity can be verified by solving a single LP (cf. Proposition 3.1). In summary, except when verifying B-stationarity, it is rarely necessary to solve the LPEC to global optimality – and even then, the task often reduces to solving a single LP.
In this paper, we focus on arguably the easiest-to-implement approach for solving LPECs, which does not suffer from the drawbacks of regularization methods. We consider the equivalent MILP formulation of LPEC (15Hu2012 :
@@
The MILP (18) can be solved with state-of-the-art commercial and open-source solvers such as Gurobi Gurobi or HiGHS Huangfu2018 . We discuss some special properties of the MILP (18), which are favorable for MILP solvers Gurobi ; Huangfu2018 . We denote the optimal objective value of (18) by $V_{\mathrm{MILP}}(x^{k})$, and of its LP relaxation, that is, a problem where $y\in\{0,1\}^{m}$ is replaced by $y\in[0,1]^{m}$, by $V_{\mathrm{LP}}(x^{k})$. Thus, it holds that $V_{\mathrm{LP}}(x^{k})\leq V_{\mathrm{MILP}}(x^{k})$. The tighter the relaxation, i.e., the smaller $|V_{\mathrm{MILP}}(x^{k})-V_{\mathrm{LP}}(x^{k})|$ is, the less branching may be needed in MILP solvers Wolsey1998 . Big-M formulations can be inefficient because they can lead to rather loose relaxations of the MILP Hu2008 . However, for $\rho<\frac{M}{2}$, the trust-region constraint (18e) always dominates the bounds on $d$ from the big-M constraints (18c) and (18d). Thus, $M$ only needs to be large enough to ensure the feasibility of (18c)–(18d). A suitable choice is $M=\max(x_{1}^{k},x_{2}^{k})+\rho$. In addition, observe that the constraint (18e) makes the feasible set of the MILP compact. Moreover, for feasible points $x^{k}$, $\rho$ can be arbitrarily small, leading to a smaller $V_{\mathrm{LP}}(x^{k})$ and thus tighter relaxations of the MILP.
Furthermore, given a feasible complementarity pair, e.g. from the previous BNLP, a feasible binary vector $y$ can be constructed, providing an upper bound for $V_{\mathrm{MILP}}(x^{k})$. The availability of an upper bound can reduce branching in branch-and-bound MILP algorithms Wolsey1998 , and our implementation includes the procedure for computing such a feasible binary vector. Our numerical experiments indicate that these factors typically result in not large numbers of LP that need to be solved when solving the MILP (18). Interestingly, in our experiments the MILP formulations for LPECs are often substantially faster than solving the LPEC (15) with noncombinatorial regularization-based methods, cf. Appendix C.
We discuss a special case that frequently arises in our numerical experiments, namely when $x^{*}\in\Omega$ is S-stationary. In this case, the MILPs are particularly easy to solve since the relaxation is tight, i.e., $V_{\mathrm{MILP}}(x^{*})=V_{\mathrm{LP}}(x^{*})$. Thus, verifying B-stationarity reduces to solving a single LP in the MILP solver. We prove this for the reduced LPEC (12). However, if $\rho<\bar{\rho}$ (with $\bar{\rho}$ from (16)) and the preprocessing step for fixing integers is applied, then the feasible sets of the full MILP (18) and reduced MILP (19) coincide.
@@
Let $x^{*}\in\Omega$ be an S-stationary point of the MPEC (1) at which the MPEC-MFCQ holds. Consider the MILP reformulation of the reduced LPEC (12):
@@
While Algorithm 1 solves NLPs and MILPs similar to many MINLP methods Belotti2013 , there are significant differences in termination criteria and the particular subproblems that are solved. Nonlinear branch-and-bound algorithms terminate when the optimality gap falls below a threshold Belotti2013 , which does not directly relate to B-stationarity unless the global optimum is found. Outer approximation algorithms also solve sequences of NLPs and MILPs Belotti2013 , but their MILPs include linearized constraints from multiple different linearization points, making them different from the MILP (18). Consequently, it is not immediately clear when these methods terminate at B-stationary points. In summary, Algorithm 1 is not a MINLP algorithm; rather, it employs MILP solvers as a tool for solving LPECs. It aims to find a locally stationary point of an MPEC, which—unlike in MINLPs – is a well-defined notion.